OpenStudy (anonymous):

prove that Square root 2 + square root 3 is irrational. step by step

OpenStudy (anonymous):

well, you are adding two irrational numbers, so why would the sum be rational? it wouldn't be.

OpenStudy (kinggeorge):

@student_basil Both \(\sqrt{2}\) and \(1-\sqrt2\) are irrational, yet their sum is 1, which is rational. So you can't simply say that \(\sqrt2+\sqrt3\) is irrational because both \(\sqrt2\) and \(\sqrt3\) are irrational. What I would do instead is suppose that \(\sqrt2+\sqrt3\) were rational. Then\[\sqrt2+\sqrt3=\frac{a}{b}\]for some integers \(a,b\). Then we square both sides to get\[5+2\sqrt6=\frac{a^2}{b^2}.\]So\[\sqrt{6}=\frac{a^2-5b^2}{2b^2}.\]This would imply that \(\sqrt6\) is rational, which is clearly a contradiction. So \(\sqrt2+\sqrt3\) is irrational.

OpenStudy (anonymous):

Maybe use a lemma to prove that \[\sqrt{6}\]is indeed irrational. Use the same technique you would use to prove any number is irrational, which is pretty much what KingGeorge used.